OpenStudy (anonymous):

4. A company manufactures and sells video games. A survey of video game stores indicated that at a price of $93 each, the demand would be 200 games, and at a price of $43 each, the demand would be 1,200 games. If a linear relationship between price and demand exists, which of the following equations models the price-demand relationship? (Let x represent the price per video game and y represent the demand.)

OpenStudy (anonymous):

Find slope m = (y2-y1) / (x2-x1) find b in the equation: y = mx + b There, you'll have your equation for the linear relationship.

OpenStudy (anonymous):

here are the choices,

OpenStudy (anonymous):

you have to do what i told you. The only way you can learn :)

OpenStudy (anonymous):

I keep getting the wrong answer on paper, and I don't have a decent calculator

OpenStudy (anonymous):

Also I have dyscalculia

OpenStudy (anonymous):

oh my bad, sorry for the late reply

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

give me a sec

OpenStudy (anonymous):

the test was restarted I have another question now ..

OpenStudy (anonymous):

oh ok